A coffee manufacturer wants to market a new blend of coffee that sells for $3.92 per pound by mixing two coffees, one that sells for $2.75 per pound, and another that sells for $5.00 per pound. What amounts of each coffee should be blended to obtain 100 pounds of the desired mixture? don't know how to write this equation
Answer by mananth(16949)

--------price ----------------quantity
coffee I 2.75----------------xlb
coffee II 5----------------100-x lb
Total 3.92% ---------------- 100
2.75x+5(100-x)= 3.92* 100
2.75x+500-5x=392
2.75x-5x=392-500
-2.25x=-108
/-2.25
x=48 lb $2.75
52 lb $5.00